The author's home (4 people) used 4 m ^3 of water over a 9 day period . What does this equate to in Litres per person per day?



1
Expert's answer
2021-01-27T09:31:32-0500

Since 4 people used 4 m^3 of water over a 9 day period, then the amount of water used by 4 people per day is:


"\dfrac{4m^3}{9\space day} = \dfrac{4}{9}\dfrac{m^3}{day}"

The amount of water used by one person per day is:


"\dfrac{4}{9}\dfrac{m^3}{day} /4\space person = \dfrac{1}{9}\dfrac{m^3}{day\cdot person}"

1 m^3 is 1000 litres, thus:


"\dfrac{1}{9}\dfrac{m^3}{day\cdot person} = \dfrac{1000}{9}\dfrac{litres}{day\cdot person} \approx 111.11\dfrac{litres}{day\cdot person}"

Answer. "111.11\dfrac{litres}{day\cdot person}".